How much did the winning team have if they scored 12more than the losing team and the sum of the game was 72
Answer by Cromlix(4381)

Winning team = x + 12: Losing team = x
x + 12 + x = 72
2x = 72 - 12
2x = 60
x = 30
Winning team scored 42 points
Losing team scored 30 points
Hope this helps
:-)